# Babak Ghanbari    Robert Šámal

### v1

@@ Abstract
We present an algorithm for finding a perfect matching in a $3$-edge-connected cubic graph that intersects every $3$-edge cut in exactly one edge. Specifically, we propose an algorithm with a time complexity of $O(n\log^{4}n)$, which significantly improves upon the previously known $O(n^{3})$-time algorithms for the same problem. The technique we use for the improvement is efficient use of cactus model of 3-edge cuts. As an application, we use our algorithm to compute embeddings of $3$-edge-connected cubic graphs with limited number of singular edges (i.e., edges that are twice in the boundary of one face) in $O(n\log^{4}n)$ time; this application contributes to the study of the well-known Cycle Double Cover conjecture.
@@ 1 Introduction
In this paper we study the complexity of finding a perfect matching $M$ in a bridgeless cubic graph such that in every 3-edge cut $M$ contains exactly one edge. By parity, $M$ can contain one or three edges in a 3-edge cut; so our condition says no 3-edge cut is contained in $M$. We shortly express this by saying $M$ is well spread.
Kaiser et al. [KKN06] showed that a wells spread perfect matching exists in every bridgeless cubic graph (and went on to use this to find how much of the graph can be covered by two, three, etc. perfect matchings). Their proof uses Edmonds perfect matching polytope (vector $(1/3,1/3,\dots)$ is a fractional perfect matching, thus it is a convex combination of perfect matchings – each of them is well spread). This, however, doesn’t yield an efficient algorithm.
In [BIT13], Boyd et al. focus on developing algorithms for finding $2$-factors in bridgeless cubic graphs that cover specific edge-cuts, which brings these $2$-factors closer to Hamiltonian cycles. They provide an efficient algorithm that finds a minimum-weight $2$-factor that covers all $3$-edge cuts in weighted bridgeless cubic graphs (in contrast with finding a Hamiltonian cycle, which is computationally hard). They also provide both a polyhedral description of such $2$-factors and of their complements – well spread perfect matchings. This is the first known polynomial-time algorithm for this problem, with a time complexity of $O(n^{3})$, where $n$ is the number of vertices. In this work, we improve this result for $3$-edge-connected cubic graphs using Algorithm 1 with time complexity $O(n\log^{4}n)$.
@@
We then proceed to apply this result to the study of the well-known Cycle Double Cover (or CDC) Conjecture [Sze73, Sey79]. In the language of graph embedding, CDC conjecture is equivalent to every bridgeless cubic graph having a surface embedding with no singular edges (i.e., with no edge that is on the boundary of one face twice). As an application of our work, for a $3$-edge-connected cubic graph we can find embeddings with a bounded number of singular edges in time $O(n\log^{4}n)$.
@@ 2 Preliminaries
Petersen’s theorem [Pet91] states that every bridgeless cubic graph contains a perfect matching. Schönberger [Sch35] proved the following strengthened form of Petersen’s theorem.
Theorem 2.1 ([Sch35])
@@ 3 The cactus model
Let $G$ be a graph. A pair $(T,\varphi)$ with $\varphi:V(G)\to V(T)$ is called a cactus representation (or cactus model) for graph $G$ if it meets the following criteria:
@@
    For an arbitrary minimum cut $\{S,V(T)-S\}\in C(T)$, the cut $\{X,\bar{X}\}$ in $G$ defined by $X=\{u\in V(G)\mid\varphi(u)\in S\}$, and $\bar{X}=\{u\in V\mid\varphi(u)\in V(T)-S\}$ is a minimum cut in $G$.
@@
    Conversely, for every minimum cut $\{X,\bar{X}\}$ in $G$, there exists a minimum cut $\{S,V(T)-S\}\in C(T)$ such that $X=\{u\in V\mid\varphi(u)\in S\}$, and $\bar{X}=\{u\in V\mid\varphi(u)\in V(T)-S\}.$
@@
It was shown in [DKL76] that every $G$ has a cactus representation using a cactus graph – a multigraph where every edge is exactly one cycle. A special case of such graph is a tree with every edge doubled. It is known that when the connectivity is odd, the cactus representation is of this special type. In this case we will just call it a tree of cuts (or cactus tree). For a $3$-edge-connected cubic graph $G$, the vertices of the graph $G$ correspond to the leaf nodes of the cactus tree $T$ and all the interior nodes of $T$ are empty nodes (see Fig. 1). We will also use the following estimate for the size of $T$.
Figure 1: A 3-edge-connected cubic graph $G$ and its cactus tree $T$. Small white circles represents empty vertices. Dashed lines represent a non-trivial $3$-edge cut in $G$ and its equivalent edge in $T$.
@@
4 Well spread perfect matchings in bridgeless cubic graphs
@@
Let $G$, $A$, $G_{1}$ and $G_{2}$ be as above. Let $(\varphi,T)$ be the cactus tree for $G$. Let $e=x_{1}x_{2}$ be the edge of $T$ corresponding to $A$; that is the components of $T-e$ are $T^{\prime}_{1}$, $T^{\prime}_{2}$, $x_{i}\in V(T^{\prime}_{i})$ and $A=\varphi^{-1}(V(T^{\prime}_{2}))$. Put $T_{i}=T^{\prime}_{i}+e$ for $i=1,2$. Define $\varphi_{1}$ as a restriction of $\varphi$ to ${\bar{A}}$ and put $\varphi_{1}(a)=x_{1}$. We define $\varphi_{2}$ symmetrically.
@@
To this end, we pick root in the cut model $T$. Let $e_{x}$ be the edge connecting $x$ to its parent. We let $C_{x}$ be the three edges in the cut corresponding to $e_{x}$. We let $A_{x}$ be the vertices of $G$ that are mapped by $\varphi$ to descendants of $x$ (including $x$), so that $C_{x}=\delta(A_{x})$.
We do not use this definition directly, instead we recursively update $A_{x}$, $C_{x}$ while traversing the tree. We can compute $A_{x}$ as a union of $A_{y}$ for all children $y$ of $x$. We can compute $C_{x}$ by going over $C_{y}$ for all children $y$ of $x$, excluding edges that lead withing $A_{x}$ (between $A_{y}$ and $A_{y^{\prime}}$ for two children of $x$). To do this efficiently we utilize the Union-Find algorithm.
In a typical step of decomposition part of the algorithm, we use Theorem 4.2, with the new graphs being $G_{x}$ (which we store for later) and updated version of $G$. To illustrate the algorithm, in Figure 2 we show one step including the updated trees. We see that $T_{x}$ is a star – which means $G_{x}$ is internally 4-edge-connected. Updated $T$ simply removes children of $x$. However, we actually do not need to update $T$ and compute $T_{x}$ in the algorithm.
In a typical step of assembling part of the algorithm, we use Theorem 4.1. For graph $G$ we have our well-spread matching from the recursion. For the other graph, $G_{x}$, we compute it easily, as the graph is internally 4-edge-connected, so any perfect matching will do. Figure 3 illustrates this part of the algorithm.
@@
The output of Algorithm 1 is a well spread perfect matching.
@@
A $3$-edge-connected cubic graph $G$. A well spread perfect matching $M$.
@@
$b\leftarrow\text{neighbor of}~{}a_{x}~{}\text{in}~{}M$
$a\leftarrow\text{neighbor of}~{}b~{}\text{in}~{}C_{x}$
@@ 4.1 Complexity Analysis
Algorithm 1 finds a well spread perfect matching $M$ in a $3$-edge-connected cubic graph $G$ in time $O(n\log^{4}n)$.
@@ 5 Application
To prove Theorem 5.1, in [GŠ24], we first construct a perfect matching $M_{1}$ such that $M_{1}$ contains no odd cut of size 3. This guarantees the existence of another perfect matching $M_{2}$ satisfying $|M_{1}\cap M_{2}|\leq\frac{n}{10}$ (see [KKN06]). Next, we show that $M_{1}\cap M_{2}$ can be extended to an embedding using Lemma 2, where the potential singular edges of the embedding are precisely those in $M_{1}\cap M_{2}$, and this completes their proof.
@@ 6 Conclusion
We studied a new approach to finding a perfect matching that includes all $3$-edge cuts in a $3$-edge-connected cubic graph, utilizing the cactus representation of the graph. We referred to such a perfect matching as well spread. In Section 4, we presented an algorithm that finds a well-spread perfect matching in a $3$-edge-connected cubic graph in time $O(n\log^{4}n)$.

### v2

@@ Abstract
We present an algorithm for finding a perfect matching in a $3$-edge-connected cubic graph that intersects every $3$-edge cut in exactly one edge. Specifically, we propose an algorithm with a time complexity of $O(n\log^{4}n)$, which significantly improves upon the previously known $O(n^{3})$-time algorithms for the same problem. The technique we use for the improvement is efficient use of the cactus model of 3-edge cuts. As an application, we use our algorithm to compute embeddings of $3$-edge-connected cubic graphs with limited number of singular edges (i.e., edges that are twice in the boundary of one face) in $O(n\log^{4}n)$ time; this application contributes to the study of the well-known Cycle Double Cover conjecture.
@@ 1 Introduction
In this paper we study the complexity of finding a perfect matching $M$ in a bridgeless cubic graph such that in every 3-edge cut $M$ contains exactly one edge. By parity, $M$ can contain one or three edges in a 3-edge cut; so our condition says no 3-edge cut is contained in $M$. We shortly express this by saying $M$ is well-spread.
Kaiser et al. [KKN06] showed that a well-spread perfect matching exists in every bridgeless cubic graph (and went on to use this to find how much of the graph can be covered by two, three, etc. perfect matchings). Their proof uses Edmonds’ perfect matching polytope (vector $(1/3,1/3,\dots)$ is a fractional perfect matching, thus it is a convex combination of perfect matchings – each of them is well-spread). This, however, doesn’t yield an efficient algorithm.
In [BIT13], Boyd et al. focus on developing algorithms for finding $2$-factors in bridgeless cubic graphs that cover specific edge-cuts, which brings these $2$-factors closer to Hamiltonian cycles. They provide an efficient algorithm that finds a minimum-weight $2$-factor that covers all $3$-edge cuts in weighted bridgeless cubic graphs (in contrast with finding a Hamiltonian cycle, which is computationally hard). They also provide both a polyhedral description of such $2$-factors and of their complements – well-spread perfect matchings. This is the first known polynomial-time algorithm for this problem, with a time complexity of $O(n^{3})$, where $n$ is the number of vertices. In this work, we improve this result for $3$-edge-connected cubic graphs using Algorithm 1 with time complexity $O(n\log^{4}n)$.
@@
We then proceed to apply this result to the study of the well-known Cycle Double Cover (or CDC) Conjecture [Sze73, Sey79]. In the language of graph embedding, the CDC conjecture is equivalent to every bridgeless cubic graph having a surface embedding with no singular edges (i.e., with no edge that is on the boundary of one face twice). As an application of our work, for a $3$-edge-connected cubic graph we can find embeddings with a bounded number of singular edges in time $O(n\log^{4}n)$.
@@ 2 Preliminaries
Petersen’s theorem [Pet91] states that every bridgeless cubic graph contains a perfect matching. Schönberger [Sch34] proved the following strengthened form of Petersen’s theorem.
Theorem 2.1 ([Sch34])
@@ 3 The Cactus Model
Let $G$ be a graph. A pair $(T,\varphi)$ with $\varphi:V(G)\to V(T)$ is called a cactus representation (or cactus model) for the graph $G$ if it meets the following criteria:
@@
    For an arbitrary minimum cut $\{S,V(T)-S\}\in C(T)$, the cut $\{X,\bar{X}\}$ in $G$ defined by $X=\{u\in V(G)\mid\varphi(u)\in S\}$, and $\bar{X}=\{u\in V(G)\mid\varphi(u)\in V(T)-S\}$ is a minimum cut in $G$.
@@
    Conversely, for every minimum cut $\{X,\bar{X}\}$ in $G$, there exists a minimum cut $\{S,V(T)-S\}\in C(T)$ such that $X=\{u\in V(G)\mid\varphi(u)\in S\}$, and $\bar{X}=\{u\in V(G)\mid\varphi(u)\in V(T)-S\}.$
@@
It was shown in [DKL76] that every $G$ has a cactus representation using a cactus graph – a multigraph where every edge is in exactly one cycle. A special case of such graph is a tree with every edge doubled. It is known that when the connectivity is odd, the cactus representation is of this special type. In this case we will just call it a tree of cuts (or cactus tree). For a $3$-edge-connected cubic graph $G$, the vertices of the graph $G$ correspond to the leaf nodes of the cactus tree $T$ and all the interior nodes of $T$ are empty nodes (see Fig. 1). We will also use the following estimate for the size of $T$.
Figure 1: A 3-edge-connected cubic graph $G$ and its cactus tree $T$. Small white circles represent empty vertices. Dashed lines represent a non-trivial $3$-edge cut in $G$ and its equivalent edge in $T$.
@@
4 Well-Spread Perfect Matchings in Bridgeless Cubic Graphs
@@
Let $G$, $A$, $G_{1}$, $G_{2}$, $A$ and ${\bar{A}}$ be as above. Let $(\varphi,T)$ be the cactus tree for $G$. Let $e=x_{1}x_{2}$ be the edge of $T$ corresponding to $A$; that is the components of $T-e$ are $T^{\prime}_{1}$, $T^{\prime}_{2}$, $x_{i}\in V(T^{\prime}_{i})$ and $A=\varphi^{-1}(V(T^{\prime}_{2}))$. Put $T_{i}=T^{\prime}_{i}+e$ for $i=1,2$. Define $\varphi_{1}$ as a restriction of $\varphi$ to ${\bar{A}}$ and put $\varphi_{1}(a)=x_{1}$. We define $\varphi_{2}$ symmetrically.
@@
To this end, we pick a root in the cut model $T$. For a vertex $x$, let $e_{x}$ be the edge connecting $x$ to its parent. We let $C_{x}$ be the three edges in the cut corresponding to $e_{x}$. We let $A_{x}$ be the vertices of $G$ that are mapped by $\varphi$ to descendants of $x$ (including $x$), so that $C_{x}=\delta(A_{x})$.
We do not use this definition directly, instead we recursively update $A_{x}$, $C_{x}$ while traversing the tree. We can compute $A_{x}$ as a union of $A_{y}$ for all children $y$ of $x$. We can compute $C_{x}$ by going over $C_{y}$ for all children $y$ of $x$, excluding edges that lie within $A_{x}$ (between $A_{y}$ and $A_{y^{\prime}}$ for two children of $x$). To do this efficiently we utilize the Union-Find algorithm.
In a typical step of the decomposition part of the algorithm, we use Theorem 4.2, with the new graphs being $G_{x}$ (which we store for later) and updated version of $G$. To illustrate the algorithm, in Figure 2 we show one step including the updated trees. We see that $T_{x}$ is a star – which means $G_{x}$ is internally 4-edge-connected. Updated $T$ simply removes children of $x$. However, we actually do not need to update $T$ and compute $T_{x}$ in the algorithm.
In a typical step of the assembling part of the algorithm, we use Theorem 4.1. For graph $G$ we have our well-spread matching from the recursion. For the other graph, $G_{x}$, we compute it easily, as the graph is internally 4-edge-connected, so any perfect matching will do. Figure 3 illustrates this part of the algorithm.
@@
The output of Algorithm 1 is a well-spread perfect matching.
@@
A $3$-edge-connected cubic graph $G$. A well-spread perfect matching $M$.
@@
$b\leftarrow\text{vertex }~{}s.t.~{}a_{x}b\in M$
$a\leftarrow\text{vertex }~{}s.t.~{}ab\in C_{x}$
@@ 4.1 Complexity Analysis
Algorithm 1 finds a well-spread perfect matching $M$ in a $3$-edge-connected cubic graph $G$ in time $O(n\log^{4}n)$.
@@ 5 Application
To prove Theorem 5.1, in [GŠ24], we first construct a perfect matching $M_{1}$ such that $M_{1}$ contains no odd cut of size 3. This guarantees the existence of another perfect matching $M_{2}$ satisfying $|M_{1}\cap M_{2}|\leq\frac{n}{10}$ (see [KKN06]). Next, decompose $G\backslash M_{1}$ and $M_{1}\Delta M_{2}$ into circuits $C_{1},C_{2},\dots C_{t}$. These circuits form a partial CDC and by Lemma 2 can be extended to an embedding in which the only potential singular edges are those that are not covered by $C_{1},C_{2},\dots C_{t}$, that is, edges in $M_{1}\cap M_{2}$. See [GŠ24] for the details.
@@ 6 Conclusion
We studied a new approach to finding a perfect matching that includes all $3$-edge cuts in a $3$-edge-connected cubic graph, utilizing the cactus representation of the graph. We referred to such a perfect matching as well-spread. In Section 4, we presented an algorithm that finds a well-spread perfect matching in a $3$-edge-connected cubic graph in time $O(n\log^{4}n)$.